OpenStudy (anonymous):

For the circle with equation (x -4)^2 + (y +2)^2 = 9

OpenStudy (anonymous):

The coordinates of the center are: Radius of the circle is: Diameter of the circle is:

OpenStudy (anonymous):

do you know the standard eqn of a circle?

OpenStudy (anonymous):

\[(x-h)^2+ (y-k)^2=r^2\]

OpenStudy (anonymous):

(x – h)^2 + (y – k)^2 = r2

OpenStudy (anonymous):



OpenStudy (anonymous):

exactly (h,k) are your coordinates, and r is your radius

OpenStudy (anonymous):

well i guess peach has solved it

OpenStudy (anonymous):

oh w8 radius is 3 diameter would be 6?

OpenStudy (anonymous):

and (4,-2) for center

OpenStudy (anonymous):

you have your signs for the coordinates mixed up. But the radius is right, 3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nvm you have it right lol

OpenStudy (anonymous):

ty ^^